OpenStudy (anonymous):

In an 8-woman race, gold, silver and bronze medals are to be given to the first, second, and third place finishers. Assuming no ties, in how many ways can these medals be awarded? A. 9 ways B. 24 ways C. 336 ways D. 40,320 ways

OpenStudy (anonymous):

@punksauce

OpenStudy (anonymous):

@.WhiteDragon.

OpenStudy (punksauce):

I'm pretty sure it would be B. but it depends on what your lesson is about.

OpenStudy (anonymous):

permutations

OpenStudy (ddcamp):

8 possible options for first place, 7 possible for second place, and 6 possible for third 8×7×6=336 possible ways.

OpenStudy (anonymous):

thanx punksauce and ddcamp I got 100%

OpenStudy (ddcamp):

Or using the combu formula: n=8 (total runners) r=3 (people getting medals) \[nPr = \frac{n!}{(n-r)!}=\frac{8\times7\times6\times5...}{6\times5\times4...} = \frac{6\times7\times6}{1}=336\]